Two point charges are placed as follows: charge q_{1} = - 1.50nC is at y = 6.00 m and charge q_{2} = 3.20nC is at the origin. What is the total force (magnitude and direction) exerted by these two charges on a negative point charge q_{3} = - 5.00nC located at (2.00 m, -4.00 m)?

1
Expert's answer
2022-03-25T15:02:12-0400

Force

"F=F_1+F_2"

"F=\frac{kq_3}{1}[\frac{q_1}{r^2}+\frac{q_2}{r^2}]"

"F=-\frac{9\times10^9 \times 5\times10^{-9}}{1}[\frac{1.50\times10^{-9}}{(\sqrt{104})^2}+\frac{3.20\times10^{-9}}{(\sqrt20)^2}]"

"F=-45\times(0.01442+0.16)\times10^{-9}=-7.8489\times10^{-9}N"